# PuzzleBench: Can LLMs Solve Challenging First-Order Combinatorial Reasoning Problems?

### v2

@@ Abstract
Recent works show that the largest of the large language models (LLMs) can solve many simple reasoning tasks expressed in natural language, without any/much supervision. But, can they also solve challenging first-order combinatorial reasoning problems, such as graph coloring, knapsack and cryptarithmetic? To answer this question, we present PuzzleBench, a dataset of $31$ such challenging problems along with a few solved instances for each problem. These problems are all first order, i.e., they can be instantiated with problem instances of varying sizes, and most of them are NP-hard, requiring several reasoning steps to reach the solution. We first observe that LLMs, even when aided by symbolic solvers, perform rather poorly on our dataset. In response, we propose a new approach, Puzzle-LM, which combines LLMs with both symbolic solvers and program interpreters, along with feedback from solved examples, to achieve huge performance gains. Our extensive experimentation and analyses offer new insights into the reasoning abilities and limitations of present-day LLMs.
PuzzleBench: Can LLMs Solve Challenging First-Order Combinatorial Reasoning Problems?
Chinmay Mittal     Krishna Kartik     Mausam     Parag Singla Indian Institute of Technology, Delhi {chinmay.mittal.iitd, kartikkrishna108}@gmail.com     {mausam, parags}@cse.iitd.ac.in
@@ 1 Introduction
Recent works have shown that large language models (LLMs) such as GPT4 OpenAI (2023) can reason like humans Wei et al. (2022a), and solve diverse natural language reasoning tasks, without the need for any fine-tuning Wei et al. (2022c); Kojima et al. (2022); Zhou et al. (2023); Zheng et al. (2023). We note that, while impressive, most of the tasks studied are simple reasoning problems, generally requiring only a handful of reasoning steps to reach a solution.
@@
We are motivated by the goal of assessing the reasoning limits of modern-day LLMs. In this paper, we study the computationally intensive, first-order combinatorial problems posed in natural language. These problems (e.g., sudoku, knapsack, graph coloring, cryptarithmetic) have long served as important testbeds to assess the intelligence of AI systems Russell and Norvig (2010), and strong traditional AI methods have been developed for them. Can LLMs solve these directly? If not, can they solve these with the help of symbolic AI systems like SMT solvers?
@@
To answer these questions, we release a dataset named PuzzleBench, consisting of $31$ such problems along with a few solved instances of each problem (see Figure 1). We refer to such problems as puzzles / fcore (first-order combinatorial reasoning) problems. Fcore problems can be instantiated with instances of varying sizes, e.g., 9$\times$9 and 16$\times$16 sudoku. Most of the problems in PuzzleBench are NP-hard and solving them will require extensive planning and search over a large number of combinations. Not surprisingly, our initial experiments reveal that the best of LLMs can only solve less than a third of these problems.
@@
We then turn to recent approaches that augment LLMs with tools for better reasoning. Program-aided Language models (PAL) Gao et al. (2023) uses LLMs to generate programs, offloading execution to a program interpreter. Logic-LM Pan et al. (2023) and SAT-LM Ye et al. (2023) use LLMs to convert questions to symbolic representations, and external symbolic solvers perform the actual reasoning. Our experiments show that, by themselves, their performances are not that strong on PuzzleBench. At the same time, both these methods demonstrate complementary strengths – PAL can handle first-order structures well, whereas Logic-LM is better at complex reasoning.
In response, we propose a new approach named Puzzle-LM, which combines the powers of both PAL and symbolic solvers with LLMs. In particular, the LLM generates an instance-agnostic program for an fcore problem that converts any problem instance to a symbolic (in our case, SMT) representation. This program passes this representation to a symbolic SMT solver, Z3 De Moura and Bjørner (2008), which returns a solution back to the program. The program then converts the symbolic solution to the desired output representation, as per the natural language instruction. Interestingly, in contrast to LLMs with symbolic solvers, once this program is generated, inference on new fcore instances can be done without any LLM calls.
@@
Puzzle-LM outperforms few-shot prompting by $29.98$, PAL by $11.97$ and Logic-LM by $23.14$ percent points on PuzzleBench, with GPT-4-Turbo as the LLM. Given the structured nature of fcore problems, we also show how solved examples of small sizes can help improve the overall reasoning. Utilizing feedback from solved examples for just four rounds yields a further $20.5$ percent points gain for Puzzle-LM. We perform additional analyses to understand the impact of other hyperparameters, and also describe the errors made by Puzzle-LM. We release the dataset and code for further research.
@@ 2 Related Work
Neuro-Symbolic AI: Our work falls in the broad category of neuro-symbolic AI Yu et al. (2023) which builds models leveraging the complementary strengths of neural and symbolic methods. Several prior works build neuro-symbolic models for solving combinatorial reasoning problems Palm et al. (2018); Wang et al. (2019); Paulus et al. (2021); Nandwani et al. (2022a, b). These develop specialized problem-specific modules (that are typically not size-invariant), which are trained over large training datasets. In contrast, Puzzle-LM uses LLMs and bypasses problem-specific architectures, generalizes to problems of varying sizes and is trained with very few solved instances.
@@
Reasoning with Language Models: The previous paradigm to reasoning was fine-tuning of LLMs Clark et al. (2021); Tafjord et al. (2021); Yang et al. (2022), but as LLMs scaled, they are found to reason well, when provided with in-context examples without any fine-tuning Brown et al. (2020); Wei et al. (2022b). Since then, many prompting approaches have been developed that leverage in-context learning. Prominent ones include Chain of Thought (CoT) prompting Wei et al. (2022c); Kojima et al. (2022), Least-To-Most prompting Zhou et al. (2023), Progressive-Hint prompting Zheng et al. (2023) and Tree-of-Thoughts (ToT) prompting Yao et al. (2023). These works generally study relatively simple reasoning tasks.
@@
Tool Augmented Language Models: Augmenting LLMs with external tools has emerged as a way to solve complex reasoning problems Schick et al. (2023); Paranjape et al. (2023). The idea is to offload a part of the task to specialized external tools, thereby reducing error rates. Program-aided Language models Gao et al. (2023) invoke a Python interpreter over a program generated by an LLM. Logic-LM Pan et al. (2023) and SAT-LM Ye et al. (2023) integrate reasoning of symbolic solvers with LLMs, which convert the natural language problem into a symbolic representation. Puzzle-LM falls in this category, but combines both PAL and symbolic solvers in a specific way to solve fcore problems.
Answer Verification: Several existing works verify the correctness of LLM’s answers via an external validation, such as via symbolic solvers Zhang and Callison-Burch (2023), or external databases Peng et al. (2023). In the same spirit, Puzzle-LM utilizes solved examples in training set for verifying LLM’s output, and provide feedback to LLMs, in case of errors.
Logical Reasoning Benchmarks: There are several natural language reasoning benchmarks in literature, such as LogiQA Liu et al. (2020) and ReClor Yu et al. (2020) for mixed reasoning, GSM8K Cobbe et al. (2021) for arithmetic reasoning, FOLIO Han et al. (2022) for first-order logic, PrOntoQA Saparov and He (2023) and ProofWriter Tafjord et al. (2021) for deductive reasoning, and AR-LSAT Zhong et al. (2021) for analytical reasoning. We propose PuzzleBench, which substantially extends the complexity of these benchmarks by investigating computationally hard, first-order combinatorial reasoning problems.
3 Problem Setup
A first-order combinatorial reasoning problem $\mathcal{P}$ has three components: a space of legal input problem instances ($\mathcal{X}$), a space of legal outputs ($\mathcal{Y}$), and a set of constraints ($\mathcal{C}$) that every input-output pair must satisfy. Example, for sudoku, input space is the space of grids with $n\times n$ cells ($n$ being a perfect square), with cells partially filled with numbers from $1$ to $n$. Output space is a fully-filled grid of the same size. And, constraints comprise row, column, and box alldiff constraints, along with input cell persistence.
To communicate a structured problem instance (or its output) to an NLP system, it must be serialized in text. We overload $\mathcal{X}$ and $\mathcal{Y}$ to denote the formats for these serialized input and output instances. Two instances of serialized inputs/outputs for sudoku are shown in Figure 2 (grey box). We are also provided (serialized) training data of input-output instance pairs, $\mathcal{D}_{\mathcal{P}}$ =$=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}$, where $x^{(i)}\in\mathcal{X},y^{(i)}\in\mathcal{Y}$, such that $(x^{(i)},y^{(i)})$ honors all constraints in $\mathcal{C}$.
Further, we verbalize all three components – input/output formats and constraints – in natural language instructions. We denote these instructions by $NL(\mathcal{X})$, $NL(\mathcal{Y})$, and $NL(\mathcal{C})$, respectively. Figure 2 illustrates these for our running example.
With this notation, we summarize our setup as follows. For an abstract fcore problem $\mathcal{P}=\langle\mathcal{X},\mathcal{Y},\mathcal{C}\rangle$, we are provided $NL(\mathcal{X})$, $NL(\mathcal{Y})$, $NL(\mathcal{C})$ and training data $\mathcal{D}_{\mathcal{P}}$, and our goal is to learn a function $\mathcal{F}$, which maps any (serialized) $x\in\mathcal{X}$ to its corresponding (serialized) solution $y\in\mathcal{Y}$ such that $(x,y)$ honors all constraints in $\mathcal{C}$.
@@ 3 Problem Setup
Figure 2: PuzzleBench Example: Filling a $n\times n$ Sudoku board along with its rules, input-output format and a couple of sample input-output pairs.
4 PuzzleBench: Dataset Construction
@@
First, we shortlist computationally challenging problems. For that we manually scan Wikipedia for diverse puzzles and NP-hard algorithmic problems. We also find puzzles from other publishing houses (e.g., Nikoli). [^1] https://www.nikoli.co.jp/en/puzzles/ From this set, we select problems (1) whose rules are easy to describe in words, (2) for whom it is easy to create a serialized input/output format, (3) for whom, the training and test datasets can be created with a reasonable programming effort, and (4) whose constraints have an SMT 2.0 encoding. This manual process led to identification of $31$ fcore problems (see Table 5 for a complete list), of which 24 are known to be NP-hard and others have unknown complexity. 10 problems are graph-based (e.g., graph coloring), 15 are grid based (e.g., sudoku), and 5 are set-based (e.g., knapsack).
We then create natural language instructions for each problem. We first develop input-output formats and create instructions for them ($NL(\mathcal{X})$, $NL(\mathcal{Y})$). Second, we rewrite all constraints/rules of the problem in our own words ($NL(\mathcal{C})$). This is done so that an LLM cannot easily invoke its prior knowledge about the same problem. For the same reason, we hide the name of the puzzle. Appendices A.2 and A.3 have detailed examples of rules and formats, respectively.
@@
Next, we create train/test data for each problem. These instances are generated synthetically when possible or else hand-crafted by human experts. A single problem instance can potentially have multiple correct solutions Nandwani et al. (2021) – all solutions are provided for each training input. The instances in the test set are typically larger in size than those in training. Because of their size, test instances may have too many solutions, and computing all of them can be expensive. Instead, we provide a verification script for each problem, which outputs the correctness of a candidate solution for any test instance. We construct 15 training instances per fcore problem, and 20-50 test instances.
5 Puzzle-LM
5.1 Preliminaries and Background
@@
Preliminaries: In the following, we assume that we have access to an LLM $\mathcal{L}$, which can work with various prompting strategies, a program interpreter $\mathcal{I}$, which can execute programs written in its programming language and a symbolic solver $\mathcal{S}$, which takes as input a pair of the form $(E,V)$, where $E$ is set of equations (constraints) specified in the language of $\mathcal{S}$, and $V$ is a set of (free) variables in $E$, and produces an assignment $\mathcal{A}$ to the variables in $V$ that satisfies the set of equations in $E$. Given the an fcore problem $\mathcal{P}=\langle\mathcal{X},\mathcal{Y},\mathcal{C}\rangle$ described by $NL(\mathcal{C})$, $NL(\mathcal{X})$, $NL(\mathcal{Y})$ and $\mathcal{D_{\mathcal{P}}}$, we would like to make effective use of $\mathcal{L}$, $\mathcal{I}$ and $\mathcal{S}$, to learn the $\mathcal{F}$, which takes any input $x\in\mathcal{X}$, and maps it to $y\in\mathcal{Y}$, such that $(x,y)$ honors the constraints in $\mathcal{C}$.
@@
  • Exclusively LLM: Many prompting strategies Wei et al. (2022c); Zhou et al. (2023); Zheng et al. (2023) make exclusive use of $\mathcal{L}$ (potentially multiple times) to represent $\mathcal{F}$, i.e., $\mathcal{L}$ is supplied with a prompt consisting of the description of $\mathcal{P}$ in the natural language, via $NL(\mathcal{C})$, $NL(\mathcal{X})$, $NL(\mathcal{Y})$, the input $x$, along with specific instructions on how to solve the problem and asked to output $y$ directly. This puts the entire burden of discovering the mapping $\mathcal{F}$ on $\mathcal{L}$, ignoring $\mathcal{S}$ and $\mathcal{I}$.
  • LLM normal-$\rightarrow$ Program: In a variation of this strategy, the LLM is prompted to instead output a program which then is interpreted by $\mathcal{I}$ on the input $x$ to produce the output $y$. This is the approach taken by methods such as PAL  Gao et al. (2023)
  • LLM + Solver: This strategy makes use of both the LLM $\mathcal{L}$ and the symbolic solver $\mathcal{S}$. Whereas the primary goal of $\mathcal{L}$ is to to act as interface for translating the problem description for $\mathcal{P}$ and the input $x$, to the language of the solver $\mathcal{S}$, and the primary burden of solving the problem is on $\mathcal{S}$, whose output is then parsed as $y$. This is the approach taken by works such as Logic-LM  Pan et al. (2023) and Sat-LM  Ye et al. (2023).

### v3

# FCoReBench: Can Large Language Models Solve Challenging First-Order Combinatorial Reasoning Problems?

@@ Abstract
Can the large language models (LLMs) solve challenging first-order combinatorial reasoning problems such as graph coloring, knapsack, and cryptarithmetic? By first-order, we mean these problems can be instantiated with potentially an infinite number of problem instances of varying sizes. They are also challenging being NP-hard and requiring several reasoning steps to reach a solution. While existing work has focused on coming up with datasets with hard benchmarks, there is limited work which exploits the first-order nature of the problem structure. To address this challenge, we present FCoReBench, a dataset of $40$ such challenging problems, along with scripts to generate problem instances of varying sizes and automatically verify and generate their solutions. We first observe that LLMs, even when aided by symbolic solvers, perform rather poorly on our dataset, being unable to leverage the underlying structure of these problems. We specifically observe a drop in performance with increasing problem size. In response, we propose a new approach, SymPro-LM, which combines LLMs with both symbolic solvers and program interpreters, along with feedback from a few solved examples, to achieve huge performance gains. Our proposed approach is robust to changes in the problem size, and has the unique characteristic of not requiring any LLM call during inference time, unlike earlier approaches. As an additional experiment, we also demonstrate SymPro-LM’s effectiveness on other logical reasoning benchmarks.
@@ 1 Introduction
Recent works have shown that large language models (LLMs) can reason like humans (Wei et al., 2022a), and solve diverse natural language reasoning tasks, without the need for any fine-tuning (Wei et al., 2022c; Zhou et al., 2023; Zheng et al., 2023). We note that, while impressive, these tasks are simple reasoning problems, generally requiring only a handful of reasoning steps to reach a solution.
@@
We are motivated by the goal of assessing the reasoning limits of modern-day LLMs. In this paper, we study computationally intensive, first-order combinatorial problems posed in natural language. These problems (e.g., sudoku, knapsack, graph coloring, cryptarithmetic) have long served as important testbeds to assess the intelligence of AI systems (Russell and Norvig, 2010), and strong traditional AI methods have been developed for them. Can LLMs solve these directly? If not, can they solve these with the help of symbolic AI systems like SMT solvers? To answer these questions, we release a dataset named FCoReBench, consisting of $40$ such problems (see Figure 1).
@@
We refer to such problems as fcore (first-order combinatorial reasoning) problems. Fcore problems can be instantiated with any number of instances of varying sizes, e.g., 9$\times$9 and 16$\times$16 sudoku. Most of the problems in FCoReBench are NP-hard and solving them will require extensive planning and search over a large number of combinations. We provide scripts to generate instances for each problem and verify/generate their solutions. Across all problems we generate 1354 test instances of varying sizes for evaluation and also provide 596 smaller sized solved instances as a training set. We present a detailed comparison with existing benchmarks in the related work (Section 2).
@@
Not surprisingly, our initial experiments reveal that even the largest LLMs can only solve less than a third of these instances. We then turn to recent approaches that augment LLMs with tools for better reasoning. Program-aided Language models (PAL) (Gao et al., 2023) use LLMs to generate programs, offloading execution to a program interpreter. Logic-LM (Pan et al., 2023) and SAT-LM (Ye et al., 2023) use LLMs to convert questions to symbolic representations, and external symbolic solvers perform the actual reasoning. Our experiments show that, by themselves, their performances are not that strong on FCoReBench. At the same time, both these methods demonstrate complementary strengths – PAL can handle first-order structures well, whereas Logic-LM is better at complex reasoning. In response, we propose a new approach named SymPro-LM, which combines the powers of both PAL and symbolic solvers with LLMs to effectively solve fcore problems. In particular, the LLM generates an instance-agnostic program for an fcore problem that converts any problem instance to a symbolic representation. This program passes this representation to a symbolic solver, which returns a solution back to the program. The program then converts the symbolic solution to the desired output representation, as per the natural language instruction. Interestingly, in contrast to LLMs with symbolic solvers, once this program is generated, inference on new fcore instances (of any size) can be done without any LLM calls.
@@
SymPro-LM outperforms few-shot prompting by $21.61$, PAL by $3.52$ and Logic-LM by $16.83$ percent points on FCoReBench, with GPT-4-Turbo as the LLM. Given the structured nature of fcore problems, we find that utilizing feedback from small sized solved examples to correct the programs generated for just four rounds yields a further $21.02$ percent points gain for SymPro-LM, compared to $12.5$ points for PAL.
We further evaluate SymPro-LM on three (non-first order) logical reasoning benchmarks from literature (Tafjord et al., 2021; bench authors, 2023; Saparov and He, 2023a). SymPro-LM consistently outperforms existing baselines by large margins on two datasets, and is competitive on the third, underscoring the value of integrating LLMs with symbolic solvers through programs. We perform additional analyses to understand impact of hyperparameters on SymPro-LM and its errors. We release the dataset and code for further research. We summarize our contributions below:
  • We formally define the task of natural language first-order combinatorial reasoning and present FCoReBench, a corresponding benchmark.
  • We provide a thorough evaluation of LLM prompting techniques for fcore problems, offering new insights into existing techniques.
  • We propose a novel approach, SymPro-LM, demonstrating its effectiveness on fcore problems as well as other datasets, along with an in-depth analysis of its performance.
Figure 1: Illustrative examples of problems in FCoReBench (represented as images for illustration).
@@ 2 Related Work
Neuro-Symbolic AI: Our work falls in the broad category of neuro-symbolic AI (Yu et al., 2023) which builds models leveraging the complementary strengths of neural and symbolic methods. Several prior works build neuro-symbolic models for solving combinatorial reasoning problems (Palm et al., 2018; Wang et al., 2019; Paulus et al., 2021; Nandwani et al., 2022a, b). These develop specialized problem-specific modules (that are typically not size-invariant), which are trained over large training datasets. In contrast, SymPro-LM uses LLMs, and bypasses problem-specific architectures, generalizes to problems of varying sizes, and is trained with very few solved instances.
@@
Reasoning with Language Models: The previous paradigm to reasoning was fine-tuning of LLMs (Clark et al., 2021; Tafjord et al., 2021; Yang et al., 2022), but as LLMs scaled, they have been found to reason well, when provided with in-context examples without any fine-tuning (Brown et al., 2020; Wei et al., 2022b). Since then, many prompting approaches have been developed that leverage in-context learning. Prominent ones include Chain of Thought (CoT) prompting (Wei et al., 2022c; Kojima et al., 2022), Least-to-Most prompting (Zhou et al., 2023), Progressive-Hint prompting (Zheng et al., 2023) and Tree-of-Thoughts (ToT) prompting (Yao et al., 2023).
@@
Tool Augmented Language Models: Augmenting LLMs with external tools has emerged as a way to solve complex reasoning problems (Schick et al., 2023; Paranjape et al., 2023). The idea is to offload a part of the task to specialized external tools, thereby reducing error rates. Program-aided Language models (Gao et al., 2023) invoke a Python interpreter over a program generated by an LLM. Logic-LM (Pan et al., 2023) and SAT-LM (Ye et al., 2023) integrate reasoning of symbolic solvers with LLMs, which convert the natural language problem into a symbolic representation. SymPro-LM falls in this category and combines LLMs with both program interpreters and symbolic solvers.
Logical Reasoning Benchmarks: There are several reasoning benchmarks in literature, such as LogiQA (Liu et al., 2020) for mixed reasoning, GSM8K (Cobbe et al., 2021) for arithmetic reasoning, FOLIO (Han et al., 2022) for first-order logic, PrOntoQA (Saparov and He, 2023b) and ProofWriter (Tafjord et al., 2021) for deductive reasoning, AR-LSAT (Zhong et al., 2021) for analytical reasoning. These dataset are not first-order i.e. each problem is accompanied with a single instance (despite the rules potentially being described in first-order logic). We propose FCoReBench, which substantially extends the complexity of these benchmarks by investigating computationally hard, first-order combinatorial reasoning problems. Among recent works, NLGraph (Wang et al., 2023) studies structured reasoning problems but is limited to graph based problems, and has only 8 problems in its dataset. On the other hand, NPHardEval (Fan et al., 2023) studies problems from the lens of computational complexity, but works with a relatively small set of 10 problems. In contrast we study the more broader area of first-order reasoning, we investigate the associated complexities of structured reasoning, and have a much large problem set (sized 40). Specifically, all the NP-Hard problems in these two datasets are also present in our benchmark.
3 Problem Setup: Natural Language First-order Combinatorial Reasoning
@@
Figure 2: FCoReBench Example: Filling a $n\times n$ Sudoku board along with its rules, input-output format, and a couple of sample input-output pairs.
A first-order combinatorial reasoning problem $\mathcal{P}$ has three components: a space of legal input instances ($\mathcal{X}$), a space of legal outputs ($\mathcal{Y}$), and a set of constraints ($\mathcal{C}$) that every input-output pair must satisfy. E.g., for sudoku, $\mathcal{X}$ is the space of partially-filled grids with $n\times n$ cells, $\mathcal{Y}$ is the space of fully-filled grids of the same size, and $\mathcal{C}$ comprises row, column, and box alldiff constraints, with input cell persistence. To communicate a structured problem instance (or its output) to an NLP system, it must be serialized in text. We overload $\mathcal{X}$ and $\mathcal{Y}$ to also denote the formats for these serialized input and output instances. Two instances for sudoku are shown in Figure 2 (grey box). We are also provided (serialized) training data of input-output instance pairs, $\mathcal{D}_{\mathcal{P}}$ =$=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}$, where $x^{(i)}\in\mathcal{X},y^{(i)}\in\mathcal{Y}$, such that $(x^{(i)},y^{(i)})$ honors all constraints in $\mathcal{C}$.
Further, we verbalize all three components – input-output formats and constraints – in natural language instructions. We denote these instructions by $NL(\mathcal{X})$, $NL(\mathcal{Y})$, and $NL(\mathcal{C})$, respectively. Figure 2 illustrates these for sudoku. With this notation, we summarize our setup as follows. For an fcore problem $\mathcal{P}=\langle\mathcal{X},\mathcal{Y},\mathcal{C}\rangle$, we are provided $NL(\mathcal{X})$, $NL(\mathcal{Y})$, $NL(\mathcal{C})$ and training data $\mathcal{D}_{\mathcal{P}}$, and our goal is to learn a function $\mathcal{F}$, which maps any (serialized) $x\in\mathcal{X}$ to its corresponding (serialized) solution $y\in\mathcal{Y}$ such that $(x,y)$ honors all constraints in $\mathcal{C}$.
4 FCoReBench: Dataset Construction
@@
First, we shortlisted computationally challenging first-order problems from various sources. We manually scanned Wikipedia  [^1] https://en.wikipedia.org/wiki/List_of_NP-complete_problems for NP-hard algorithmic problems and logical-puzzles. We also took challenging logical-puzzles from other publishing houses (e.g., Nikoli),2 and real world problems from the operations research community and the industrial track of the annual SAT competition 222 https://www.nikoli.co.jp/en/puzzles/, https://satcompetition.github.io/. From this set, we selected problems (1) that can be described in natural language (we remove problems where some rules are inherently visual), and (2) for whom, the training and test datasets can be created with a reasonable programming effort. This led to $40$ fcore problems (see Table 7 for a complete list), of which 30 are known to be NP-hard and others have unknown complexity. 10 problems are graph-based (e.g., graph coloring), 18 are grid based (e.g., sudoku), 5 are set-based (e.g., knapsack), 5 are real-world settings (e.g. car sequencing) and 2 are miscellaneous (e.g., cryptarithmetic).
Two authors of the paper having formal background in automated reasoning and logic then created the natural language instructions and the input-output format for each problem. First, for each problem one author created the input-output formats and the instructions for them ($NL(\mathcal{X})$, $NL(\mathcal{Y})$). Second, the same author then created the natural language rules ($NL(\mathcal{C})$) by referring to the respective sources and re-writing the rules. These rules were verified by the other author making sure that they were correct i.e. the meaning of the problem did not change and they were unambiguous. The rules were re-written to ensure that an LLM cannot easily invoke its prior knowledge about the same problem. For the same reason, the name of the problem was hidden.
In the case of errors in the natural language descriptions, feedback was given to the author who wrote the descriptions to correct them. In our case typically there were no corrections required except 3 problems where the descriptions were corrected within a single round of feedback. A third independent annotator was employed who was tasked with reading the natural language descriptions and solving the input instances in the training set. The solutions were then verified to make sure that the rules were written and comprehensible by a human correctly. The annotator was able to solve all instances correctly highlighting that the descriptions were correct. The guidelines utilized to re-write the rules from their respective sources were to use crisp and concise English without utilizing technical jargon and avoiding ambiguities. The rules were intended to be understood by any person with a reasonable comprehension of the language and did not contain any formal specifications or mathematical formulas. Appendices A.2 and A.3 have detailed examples of rules and formats, respectively.
Next, we created train/test data for each problem. These instances are generated programmatically by scripts written by the authors. For each problem, one author also wrote a solver and a verification script, and the other verified that these scripts and suggested corrections if needed. In all but one case the other author found the scripts to be correct. These scripts (after correction) were also verified through manually curated test cases. These scripts were then used to ensure the feasibility of instances.
@@
Since a single problem instance can potentially have multiple correct solutions (Nandwani et al., 2021) – all solutions are provided for each training input. The instances in the test set are typically larger in size than those in training. Because of their size, test instances may have too many solutions, and computing all of them can be expensive. Instead, the verification script can be used, which outputs the correctness of a candidate solution for any test instance. The scripts are a part of the dataset and can be used to generate any number of instances of varying complexity for each problem to easily extend the dataset. Keeping the prohibitive experimentation costs with LLMs in mind, we generate around 15 training instances and around 34 test instances on average per problem. In total FCoReBench has 596 training instances and 1354 test instances.
5 SymPro-LM
@@
Preliminaries: In the following, we assume that we have access to an LLM $\mathcal{L}$, which can work with various prompting strategies, a program interpreter $\mathcal{I}$, which can execute programs written in its language and a symbolic solver $\mathcal{S}$, which takes as input a pair of the form $(E,V)$, where $E$ is set of equations (constraints) specified in the language of $\mathcal{S}$, and $V$ is a set of (free) variables in $E$, and produces an assignment $\mathcal{A}$ to the variables in $V$ that satisfies the set of equations in $E$. Given the an fcore problem $\mathcal{P}=\langle\mathcal{X},\mathcal{Y},\mathcal{C}\rangle$ described by $NL(\mathcal{C})$, $NL(\mathcal{X})$, $NL(\mathcal{Y})$ and $\mathcal{D_{\mathcal{P}}}$, we would like to make effective use of $\mathcal{L}$, $\mathcal{I}$ and $\mathcal{S}$, to learn the mapping $\mathcal{F}$, which takes any input $x\in\mathcal{X}$, and maps it to $y\in\mathcal{Y}$, such that $(x,y)$ honors the constraints in $\mathcal{C}$.
@@
  • Exclusively LLM: Many prompting strategies (Wei et al., 2022c; Zhou et al., 2023) make exclusive use of $\mathcal{L}$ to represent $\mathcal{F}$. $\mathcal{L}$ is supplied with a prompt consisting of the description of $\mathcal{P}$ via $NL(\mathcal{C})$, $NL(\mathcal{X})$, $NL(\mathcal{Y})$, the input $x$, along with specific instructions on how to solve the problem and asked to output $y$ directly. This puts the entire burden of discovering $\mathcal{F}$ on the LLM.
  • LLM $\rightarrow$ Program: In strategies such as PAL  (Gao et al., 2023), the LLM is prompted to output a program, which then is interpreted by $\mathcal{I}$ on the input $x$, to produce the output $y$.
  • LLM + Solver: Strategies such as Logic-LM  (Pan et al., 2023) and Sat-LM  (Ye et al., 2023) make use of both the LLM $\mathcal{L}$ and the symbolic solver $\mathcal{S}$. The primary goal of $\mathcal{L}$ is to to act as an interface for translating the problem description for $\mathcal{P}$ and the input $x$, to the language of the solver $\mathcal{S}$. The primary burden of solving the problem is on $\mathcal{S}$, whose output is then parsed as $y$.